OpenStudy (anonymous):

math 50 help please; anyone out there understand word problems well, its dealing with polynomials

OpenStudy (apoorvk):

Yes, please post you problems, I am sure someone here can help you out.

OpenStudy (lgbasallote):

i am kinda good in algebra...might be of service...

OpenStudy (anonymous):

three tired hikers camped overnight. all they had to eat was a bag of apples. during the night, one awoke and ate one-third of the apples. later a second awoke and at one-third of the apples that remained. Much later, the third camper awoke and ate one-third of those apples yet remaining after the other two had eaten. when they got up the next morning, eight apples were left. how many apples did they begin with?

OpenStudy (apoorvk):

let initial no. of apples be x. so after first ate, apples left = x - x/3 = t second one ate 1/3rd of what remained, so, apples left after he had eaten = t - t/3 =(x - x/3) - (x - x/3)/3 = w after the third ate, remaining apples = w-w/3 =[(x - x/3) - (x - x/3)/3] - [(x - x/3) - (x - x/3)/3]/3 simplifies to: 8x/27 so this is the number of apples left finally, which is stated that it equals '8' so, 8x/27 = 8 solve for 'x', and enjoy!! :))

OpenStudy (anonymous):

oh my looks very complicated, but thank you, I hope I can figure out what it all means

OpenStudy (anonymous):

1 - 1/3 = 2/3 2/3 - 1/3 ( 2/3) = 4/9 4/9 - 1/3 ( 4/9) = 8/27 Thus 8x/ 27 = 8 => x = ...

OpenStudy (apoorvk):

yeah, it just LOOKS complicated, but it isn't at all. Just go through the equations very carefully, and you 'll strike gold!

OpenStudy (anonymous):

thank you everyone for your help

OpenStudy (anonymous):

can someone please hoelp explain to me how to do x to the fifth power plus one over x plus one? its dividing by a binomial

OpenStudy (apoorvk):

i am not sure what you meant. this?

OpenStudy (anonymous):

yes, thats cool i didnt know you could do that

OpenStudy (anonymous):

I have the answer in the book i just have no clue how they got to that

OpenStudy (apoorvk):

you can actually use the equation editor too. :) so I don't understand what you want me to do. solve that or is it already the final answer. because it is in its factorised form right now.

OpenStudy (anonymous):

I have to divide it, the answer given in the back of the book is X to the fourth power minus X to the third power plus X to the second power minus X plus one

OpenStudy (anonymous):